Please show work
2. What is the wind speed for 150 miles per hour (mph) expressed in m/s? Show your work. er 3. Figure out the SI unit of the fluid viscosity u by using the equation τ=μ-u and the fluid kinematic viscosity vt 2 dy
Answer #1

2. As we know, 2.237 mi/hr = 1m/s

So, 150 mi/hr = 150 / 2.237 m/s = 67.056 m/s

or 150 mi/hr * (m/s) / (2.237 mi/hr) = 67.056 m/s

3. As we know, \tau = \mu dU /dy

Where \tau is the stear stress = Force/ Area = Force / (Length)2

and dU / dy = Change of velocity / change of distance = (length/ time) *(1/length) = 1/time

Now, \mu = \tau * ( dy/ dU) = (Force / (Length)2) / (1/time) = Force * time / (Length)2 = N.s/m2

Again, \nu = \mu / \rho = (N.s/m2) / kg/m3 = ( kg/m.s) / ( kg/m3) = m2/s